OpenStudy (anonymous):

Given the following functions f(x) and g(x), solve (f ⋅ g)(2) and select the correct answer below: f(x) = 2x2 − 10 g(x) = x + 9 −22 22 −66 66

OpenStudy (anonymous):

@Haseeb96

OpenStudy (haseeb96):

just multiply both the functions and then put 2 at the place of x and get your answer

OpenStudy (anonymous):

im confused

OpenStudy (haseeb96):

(f . g)(x) = 2x^3 +18x^2-10 x -90 this is the term you will get after multiplication

OpenStudy (haseeb96):

now just put x=2 at the place of x and get your answer

OpenStudy (haseeb96):

(f . g) (2) = 2(2)^3 +18(2)^2-10(2) -90 now solve it get your answer I cant solve it further

OpenStudy (haseeb96):

yes @drkid15 solve it ?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

can u help

OpenStudy (haseeb96):

in what ?

OpenStudy (anonymous):

solving it

OpenStudy (haseeb96):

(f . g)(2) = 16 +72 -20 - 90 = 88 -110 = -22 is your answer

OpenStudy (anonymous):

ohhhhhh ok!